OpenStudy (unklerhaukus):

Can someone please check my work \[x^2y^{\prime\prime}-3xy^\prime+3y=0;\qquad y(1)=5;\qquad y(1)^{\prime\prime}=12\]

OpenStudy (unklerhaukus):

\[x^2y^{\prime\prime}-3xy^\prime+3y=0;\qquad y(1)=5;\qquad y(1)^{\prime\prime}=12\] \[\text{let } \ln x=t\]\[\qquad x=e^t\] \[y^{\prime\prime}+(-3-1)y^\prime+3y=0\] \[y^{\prime\prime}-4y^\prime+3y=0\]\[m^2-4m+3=0\]\[(m-1)(m-3)=0\]\[m=1,3\] \[y(t)=Ae^t+Be^{3t}\] \[y(x)=Ax+Bx^3\]\[y^\prime(x)=x+3Bx^2\]\[y^{\prime\prime}(x)=6Bx\]\[y^{\prime\prime}(1)=6B=12;\qquad B=2\]\[y(x)=Ax+2x^3\]\[y(1)=A+2=5;\qquad A=3\] \[y(x)=3x+2x^3\]

OpenStudy (valpey):

Yes.

OpenStudy (unklerhaukus):

YES

OpenStudy (valpey):

But you may want to explain your leap from \[\text{let }\ln x=t\ \ x=e^t\]to\[y''+(−3−1)y'+3y=0\]

OpenStudy (unklerhaukus):

can you help with that im not really sure of the details

OpenStudy (unklerhaukus):

\[x=e^t\]\[\frac{\text dx}{\text dt}=e^t\]\[\text dx=e^t\text dt\]